# Designing Informative Metrics for Few-Shot Example Selection

### 06 Mar 2024 v1

@@ 1 Introduction
We conduct our experiments on named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. We also devise a new dataset for privacy policies involving tagging of contextual integrity parameters (Shvartzshnaider et al., 2019). We show that CP retrieval improves accuracy significantly in all cases.
@@ 4.1 Tasks
We perform our experiments on 4 different sequence tagging tasks namely NER, POS tagging, sentence chunking and on a custom dataset of privacy policies for tagging contextual integrity (CI) parameters (Shvartzshanider et al., 2023). Each task involves the annotation of tokens in a sequence with specific labels, facilitating the extraction of valuable information and enhancing language understanding. For all datasets, we use a random set of 1000 test samples. For the CI task, we use 89 samples.
@@
4. CI parameters: We additionally created a custom dataset (which we will release) that involves the sequence tagging task of CI parameters for privacy policies which is taken from Shvartzshanider et al. (2023). CI parameters sequence tagging is a crucial task within the realm of privacy-aware NLP (Shvartzshnaider et al., 2019).
@@ 4.2 Models
We perform our experiments on the GPT-Neo series of models which has parameter counts from 125 million to 20 billion (GPT-Neo-125M,GPT-Neo-1.3B, GPT-Neo-2.7B, GPT-j-6B, GPT-NeoX-20B) and additionally 2 black box models (Davinci-002 and GPT-4) for the sequence tagging tasks of NER, POS tagging and sentence chunking. For our custom dataset for CI parameters tagging, we use Llama2 variants.
@@
Davinci-002 (Black Box) $20.30$ 40.03 $35.82$ 51.79 46.0446.0446.0446.04333Accuracy has changed with respect to the original paper for this model probably because of the the reasoning in Chen et al. (2023) 52.20
@@
CI tagging (Acc.)
Baseline CP retrieval
Llama2-13B $37.56$ 56.79
Llama2-70B $43.90$ 59.35
Table 2: Results of using complexity score for retrieval (CP retrieval) on the CI parameters tagging task using Llama2 models. Best values are in boldface.
@@ 5 Results
Our results show that CP retrieval significantly improves few-shot accuracy over the baseline across all models and considered tasks. Table 1 compares the F1 and Accuracy scores (for NER, chunking and POS tagging) for different models while keeping the number of examples constant. Table 2 provides the same for the CI tagging task on Llama models.
@@ 7 Limitations
Our methodology focuses on sequence tagging tasks and cannot be applied to tasks outside this domain such as question answering. This is because the normalized entropy metric is specifically designed for encouraging selection of sentences with a wider label variety on the token level; this helps increase the sampling of underrepresented labels. Further, we have conducted all experiments for the sequence tagging tasks in only English language. Performance with other languages has not been analyzed.

### 19 Mar 2024 v2

@@ 1 Introduction
We conduct our experiments on named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. We show that CP retrieval improves accuracy significantly in all cases.
@@ 4.1 Tasks
We perform our experiments on 3 different sequence tagging tasks namely NER, POS tagging, and sentence chunking. Each task involves the annotation of tokens in a sequence with specific labels, facilitating the extraction of valuable information and enhancing language understanding. For all datasets, we use a random set of 1000 test samples.
@@ 4.2 Models
We perform our experiments on the GPT-Neo series of models which has parameter counts from 125 million to 20 billion (GPT-Neo-125M,GPT-Neo-1.3B, GPT-Neo-2.7B, GPT-j-6B, GPT-NeoX-20B) and additionally 2 black box models (Davinci-002 and GPT-4) for the sequence tagging tasks of NER, POS tagging and sentence chunking.
@@
Davinci-002 (Black Box) $20.30$ 40.03 $35.82$ 51.79 46.0446.0446.0446.04333Accuracy has changed with respect to the original paper for this model probably because of the the reasoning in chen2023chatgpts 52.20
@@ 5 Results
Our results show that CP retrieval significantly improves few-shot accuracy over the baseline across all models and considered tasks. Table 1 compares the F1 and Accuracy scores (for NER, chunking and POS tagging) for different models while keeping the number of examples constant.
@@ 7 Limitations
Our methodology focuses on sequence tagging tasks and cannot be applied to tasks outside this domain such as question answering. This is because the normalized entropy metric is specifically designed for encouraging selection of sentences with a wider label variety on the token level; this helps increase the sampling of underrepresented labels. Further, we have conducted all experiments for the sequence tagging tasks in only the English language. Performance with other languages has not been analyzed.
Errata
The previous version of this paper contained results pertaining to a dataset involving annotating statements with contextual integrity parameters. This dataset can not be publicly shared, and all related results have been removed (since they will not be reproducible).